Amazon’s new warehouse in Boston, MA is using robots to pick the ordered goods from the shelfs and package them to be sent for the final destination. According to design requirements, one robot is supposed to handle 160 orders per hour. However, due to the complexity of order retrieval, robots have to travel long distances in the warehouse and are observed to handle 140 orders per hour. Furthermore, due to maintenance and battery charging requirements, each robot operates at 95% capacity efficiency, 365 days out of the year for 23 hours every day. What is the capacity utilization of each robot?

85.125 %

82.250 %

84.215 %

83.125 %
Answer #1

Answer: design capacity= 160 per hour

Effective capacity= 140 per hour

Efficiency= 95%= 0.95

Actual capacity= Efficiency*effective capacity= 0.95*140= 133

Utilization= actual capacity/ design capacity = 133/160= 0.83125, 83.125%

Please ask, if you have any doubts through the comment section. Do rate the answer